### v1

# Maximum Length RLL Sequences in de Bruijn Graph

Abstract
A timing and synchronization system based on a de Bruijn sequence has been proposed and studied recently for a channel associated with quantum communication that requires reliable synchronization. To avoid a long period of no-pulse in such a system on-off pulses are used to simulate a zero and on-on pulses are used to simulate a one. But, these sequences have high redundancy. To reduce the redundancy, run-length limited sequences in the de Bruijn graph are proposed for the same purpose. The maximum length of such sequences in the de Bruijn graph is studied and an efficient algorithm to construct a large set of these sequences is presented. A maximum length sequence for which the position of each window can be computed efficiently is constructed. Finally, an enumeration of the number of such sequences is given and some generalizations are discussed.
Index Terms: de Bruijn graph, necklaces, RLL sequences.
1 Introduction
In a free-space quantum key distribution, one of the goals is to synchronize the transmitted and received signals accurately. There are some efforts in designing efficient and reliable timing and synchronization systems [6, 18]. In [26] a de Bruijn sequence-based timing and synchronization system is introduced using a beacon with an on-off model. In this model, a sequence of beacon pulses is used to represent a binary de Bruijn sequence. Hence, once a sub-string of beacon pulses is received, its position is also determined uniquely. Furthermore, to consider the timing jitter performance, a long period of no-pulses is forbidden. Assume one pulse slot is used to represent a binary bit. If on-pulse is one and off-pulse is zero, then a long run of zeros in the sequence, which is a long period of no-pulses, would impact the timing jitter. In [26], two pulse slots are suggested to represent a single bit, on-on is a one and on-off is a zero so that two consecutive no-pulses are avoided. However, this scheme requires $2N$ pulse slots to represent a sequence of length $N$ in the order $n$ de Bruijn graph and it is required to receive a sub-sequence of $2n$ pulse slots to locate its position, i.e., if the sequence is long of about $2N=2^{n+1}$ bits (representing a sequence of length $N$ in the de Bruijn graph), about $2\log N$ pulse slots are required to locate its position instead of $\log N$ pulse slots if any sequence of $n$-pulses was permitted (all logarithms in this paper are in base 2). A second disadvantage in the proposed scheme is that the sequences of consecutive zeros were constrained to length one, while in reality, a few consecutive no-pulses are permitted subject to a constraint that their length will not be larger than a certain threshold $s$. Therefore, there is a target to use less redundant pulse slots to achieve both goals, to synchronize accurately, and to avoid long periods of no-pulses. For this purpose run-length limited (RLL) de Bruijn sequences are proposed [3, 4], where one pulse is represented by one binary bit. An on-pulse is represented by a one and an off-pulse is represented by a zero. This scheme is more general and more efficient with lower redundancy and higher rate than the ones in the previous work. The scheme combines two concepts, RLL sequences and sequences in the de Bruijn graph which will be defined now.
RLL sequences are binary sequences in which there is an upper bound on the number of consecutive zeros in a sequence. An $(n,s)$-word is a binary word of length $n$ in which the longest run of consecutive zeros is of length at most $s$. An $(n,s)$-sequence is a sequence whose windows of length $n$ are distinct $(n,s)$-words. The family of sequences in which there are no runs of more than $s$ consecutive zeros was extensively studied due to many applications that require such sequences [16].
The de Bruijn graph of order $n$, $G_{n}$, was defined first by Nicolaas Govert de Bruijn [1] and in parallel by Good [14]. The graph is a directed graph with $2^{n}$ vertices which are represented by the set of all binary words of length $n$. The edges of $G_{n}$ are represented by the $2^{n+1}$ binary words of length $n+1$. There is a directed edge $(x_{0},x_{1},\ldots,x_{n-1},x_{n})$ from the vertex $(x_{0},x_{1},\ldots,x_{n-1})$ to the vertex $(x_{1},\ldots,x_{n-1},x_{n})$. A span $n$ de Bruijn sequence is a cyclic binary sequence in which each binary $n$-tuple is contained exactly once in a window of length $n$. A walk (tour) in the graph is a sequence of directed edges such that the end-vertex of one edge is the start-vertex of the next edge. In a tour, the first vertex is also the last one. Any binary sequence can be represented by a walk in $G_{n}$, where any $n$ consecutive symbols represent a vertex and any $n+1$ consecutive symbols represent an edge. A cyclic sequence can be represented by a tour. The consecutive $n$ symbols and $n+1$ symbols represent a walk with its consecutive vertices and consecutive edges, respectively. A path (cycle) in the graph is a walk (tour) with no repeated vertices. In a cycle, each vertex can be considered the first vertex of the cycle. A trail (circuit) in the graph is a walk (tour) with no repeated edges. A span $n$ de Bruijn sequence can be represented by an Eulerian circuit in $G_{n-1}$, i.e., a circuit which traverses each edge exactly once. In a circuit, the first vertex is also the last one. It can be also represented by a Hamiltonian cycle in $G_{n}$, i.e., a cycle that visit each vertex of $G_{n}$ exactly once. Similarly, any path in $G_{n}$ can be represented by a sequence with no repeated $n$-tuples. This sequence can be represented also by a trail in $G_{n-1}$.
One structure that will be used in our exposition and is extensively studied in the literature associated with the de Bruijn graph is a necklace. A necklace of order $n$ is a cycle in $G_{n}$ whose length is a divisor of $n$. If the cycle is of length $n$, then the necklace is of full-order. If the necklace is of length which is a divisor of $n$ that is smaller than $n$, then the necklace is degenerated. Each vertex in $G_{n}$ is represented by a word of length $n$ and in the necklace, we have all cyclic shifts of such a word. The necklaces of order $n$ are the equivalence classes of the relation defined on words of length $n$, where two words are related if one is a cyclic shift of the other. If the necklace has length $d$ which is a divisor of $n$, then it contains $d$ words of length $n$, but the necklace can be represented by a word of length $d$. In $G_{n-1}$ such a necklace of order $n$ is also a cycle of length which is a divisor of $n$ with no repeated edges. The necklaces in $G_{n-1}$ is represented by the edges whose representation are words of length $n$. The runs of zeros in a necklace are considered to be a cyclic run since the necklace is a cycle. Let $(n,s)$-necklace be a necklace of order $n$ that does not have a (cyclic) run of more than $s$ consecutive zeros. Such a necklace contains only $(n,s)$-words.
Example 1.
Assume $n=6$ and consider the $(6,2)$-necklace with the vertex $(001011)$. This necklace is of full-order and it contains the six words of length 6 in the cycle
$(001011)\rightarrow(010110)\rightarrow(101100)\rightarrow(011001)\rightarrow(110010)\rightarrow(100101)\rightarrow(001011)$
in $G_{6}$. These six words are cyclic shifts of each other and their necklace is represented for example by $[001011]$.
In $G_{5}$ this necklace is represented by the cycle
$(00101)\rightarrow(01011)\rightarrow(10110)\rightarrow(01100)\rightarrow(11001)\rightarrow(10010)\rightarrow(00101)$
where the representation of the edges in $G_{5}$ is the same as the representation of the vertices in $G_{6}$ for these two cycles in $G_{5}$ and $G_{6}$, respectively.
The degenerated necklace $[011]$ of order 6 contains the three vertices in $G_{6}$ which form the cycle of length 3 whose vertices and edge are given by
$(011011)\rightarrow(110110)\rightarrow(101101)\rightarrow(011011)~{}.$
In $G_{5}$ this cycle is
$(01101)\rightarrow(11011)\rightarrow(10110)\rightarrow(01101)~{}.$
$\blacksquare$
In this work, we will consider a combination of sequences in the de Bruijn graph that have a run-length constraint on the number of consecutive zeros. We will be interested in the maximum length cyclic and acyclic $(n,s)$-sequences. By their definition, if $s$ is small, these sequences can be used in free-space quantum key distribution with flexibility in the parameters.
Let $G_{n}(s)$ be the subgraph of $G_{n}$ induced by all the vertices of $G_{n}$ whose representations are $(n,s)$-words. Our goal is to find the maximum length path and maximum length cycle in $G_{n}(s)$. The definition of $G_{n}(s)$ implies that the edges of this graph also do not have a run of zeros whose length is greater than $s$. This property will be important in $G_{n-1}(s)$. An upper bound on the maximum length of such a sequence will use a trail in $G_{n-1}(s)$ (which is a circuit if the sequence is cyclic) and a lower bound on the maximum length will use a path in $G_{n}(s)$ (which is a cycle if the sequence is cyclic). This is unique to our paper as usually it is not required to use different orders of the graph for similar tasks. Note, that a cyclic sequence of length $N$ with no repeated $n$-tuples has $N$ binary digits, $N$ vertices, and $N$ edges in $G_{n}$ and in $G_{n}(s)$. An acyclic sequence with no repeated $n$-tuples has $N+n-1$ binary digits, $N$ vertices, and $N-1$ edges in $G_{n}$ and $G_{n}(s)$. Henceforth, unless stated otherwise, all the $(n,s)$-sequences that will be considered will be cyclic.
Example 2.
The following cycle in $G_{6}(2)$
$\rightarrow(100100)\rightarrow(001001)\rightarrow(010011)\rightarrow(100111)\rightarrow(001110)\rightarrow(011101)\rightarrow(111010)$
$\rightarrow(110100)\rightarrow(101001)\rightarrow(010010)\rightarrow(100101)\rightarrow(001011)\rightarrow(010111)\rightarrow(101110)$
$\rightarrow(011100)\rightarrow(111001)\rightarrow(110010)$
form the $(6,2)$-sequence
$10010011101001011$
of length 17. The sequence contains two words $(001110)$ and $(011100)$ which are $(6,2)$-words not in a $(6,2)$-necklace since their necklace contains the words $(000111)$, $(100011)$, $(110001)$, and $(111000)$, that are not $(6,2)$-words.
$\blacksquare$
Example 2 indicates that there could be long $(n,s)$-sequences which contain $(n,s)$-words that are not contained in $(n,s)$-necklaces. This implies that it is not straightforward to know the maximum length of an $(n,s)$-sequence.
The rest of the paper is organized as follows. In Section 2 we consider an upper bound on the maximum length of $(n,s)$-sequences which are cycles in $G_{n}$. We prove that this bound is equal to the number of words in the $(n,s)$-necklaces. The trail with the maximum length can have another $s$ edges (and $s$ vertices) from $G_{n-1}(s)$. The proof will be based on the necklaces obtained by the edges of $G_{n-1}(s)$. In Section 3 it will be proved that the upper bound which was derived in Section 2 can be attained by many $(n,s)$-sequences. This is proved by considering Hamiltonian cycles in a subgraph of $G_{n}(s)$ that contains only the vertices of $(n,s)$-necklaces. By adding $s$ vertices (and $s$ edges) of $G_{n-1}(s)$ from necklaces which contain a cyclic run with $s+1$ zeros a longer acyclic $(n,s)$-sequence is obtained. An efficient algorithm to construct these sequences will be presented. In Section 4 enumeration of the number of $(n,s)$-necklaces is done and a formula for the length of a maximum length $(n,s)$-sequence is derived. In Section 5 a maximum length $(n,s)$-sequence from which the position of each $n$-tuple can be efficiently decoded is constructed. Conclusion, several possible generalizations, and future research are discussed in Section 6.
2 The Maximum Length of an $(n,s)$-Sequence
In this section, we will present an upper bound on the length of $(n,s)$-sequences. Let ${\cal C}$ be a circuit of maximum length in $G_{n-1}$ which does not contain an edge with $s+1$ consecutive zeros in its representation. This circuit is a cycle in $G_{n}$ with no vertex with $s+1$ consecutive zeros in its representation and hence it represents an $(n,s)$-sequence. To find an upper bound on the length of ${\cal C}$ we will try to remove the edges from $G_{n-1}(s)$ which are not contained in ${\cal C}$. These edges are associated with vertices whose in-degree is not equal to their out-degree. The following lemmas are immediate consequences from the binary representation of the vertices, their in-edges, and out-edges.
Lemma 2.1.
The vertices with in-degree one in $G_{n-1}(s)$ are all the vertices that start in their binary representation with exactly $s$ zeros followed by a one.
Lemma 2.2.
The vertices with out-degree one in $G_{n-1}(s)$ are all the vertices that end in their binary representation with a one followed by exactly $s$ zeros.
The following lemmas are simple observations from the definitions.
Lemma 2.3.
All the words in a necklace of order $n$ with at least two disjoint runs of $s+1$ or more zeros have at least one run with $s+1$ consecutive zeros.
Corollary 2.4.
All the words of each necklace of order $n$ with at least two disjoint runs of $s+1$ zeros are not contained in the circuit ${\cal C}$.
Lemma 2.5.
If a necklace of order $n$ has a word with a run of at least $12s+$ 1 zeros, then each word on this necklace has at least one run with more than $s$ zeros.
Corollary 2.6.
All the words of each necklace of order $n$ with a run of at least $12s+$ 1 zeros are not contained in the circuit ${\cal C}$.
Henceforth, a string which starts and ends with a one and has no run of more than $s$ zeros will be called an $s$-ones string.
Lemma 2.7.
In $G_{n-1}$ a necklace of order $n$ that contains a unique run with more than $s$ zeros and the length of this run is $s+k$, $1\leq k\leq s$, contains $n-1-s+k$ edges with a run of more than $s$ zeros.
Proof.
Consider the edge $0^{k-1}X0^{s+1}$, $1\leq k\leq s$, where $X$ is an $s$-ones string of length $n-s-k$. This is the first edge in $G_{n-1}$ that has a run of $s+k$ zeros, associated with a necklace of order $n$, to be considered. The last edge with a run of more than $s$ zeros on this necklace to be considered is $0^{s+1}X0^{k-1}$. Now, in the path from the edge $0^{k-1}X0^{s+1}$ to the edge $0^{s+k}X$ in this necklace there are $1+k-1+n-s-k=n-s$ edges including the first edge and the last edge. All these edges have a run of zeros with a length greater than $s$. From the edge $0^{s+k}X$ to the edge $0^{s+1}X0^{k-1}$ in the necklace there are exactly $k-1$ edges including the last edge, excluding the first edge. Again, all these edges have a run of zeros with a length greater than $s$. Hence, in a necklace with a unique long run of more than $s$ zeros whose length is exactly $s+k$, the total number of edges with a run of at least $s+1$ zeros is $n-s+k-1$. ∎
Lemma 2.8.
Any edge in $G_{n-1}(s)$ of the form $e=0^{s}X0^{k}$, where $1\leq k\leq s$, and $X$ is an $s$-ones string of length $n-s-k$, yields a path, which starts at the vertex $(0^{s}X0^{k-1})$, whose length is at least $s-k+1$, its edges do not have a run of more than $s$ zeros, and these edges are not contained in ${\cal C}$.
Proof.
The in-vertex of the edge $e=(v,u)$ has the form $v=(0^{s}X0^{k-1})$. By Lemma 2.1, the in-degree of the vertex $v$ is one and hence any circuit of $G_{n-1}(s)$ can contain at most one of its two out-edges. Remove one of its out-edges $v\rightarrow v_{1}$ that is not contained in ${\cal C}$ from $G_{n-1}(s)$. Now, $v_{1}$ has in-degree one and hence ${\cal C}$ contains at most one of its out-edges. Remove the edge which is not contained in ${\cal C}$ from $G_{n-1}(s)$. The process continues until the out-degree of the vertex that is reached is already one. The smallest number of edges removed by this process is $s-k+1$ since the shortest path from $v$ to such a vertex $(0^{k-1}X0^{s})$ is of length $s-k+1$ (by Lemma 2.2 the vertex $(0^{k-1}X0^{s})$ has out-degree one.). ∎
For a vertex of the form $v=(0^{s}X0^{k-1})$, where $1\leq k\leq s$ and $X$ is a $s$-ones string, in $G_{n-1}(s)$, let ${\cal P}(v)$ denote the sub-path of length at least $s-k+1$ in $G_{n-1}(s)$, that starts in the vertex $v$ and ends at a vertex $v^{\prime}$ whose out-degree one (as described in the proof of Lemma 2.8) and whose edges are not contained in ${\cal C}$. The sub-path ${\cal P}(v)$ will be called the deleted path of $v$.
Corollary 2.9.
After all the deleted paths induced by the proof of Lemma 2.8 are removed from $G_{n-1}(s)$ the graph remains only with vertices whose in-degree equals to their out-degree.
Proof.
This is an immediate consequence of the proof of Lemma 2.8. It should be noted that for a vertex $v$ with two out-edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$, one vertex (either $v_{1}$ or $v_{2}$) ends with a zero and hence the edge $e$ in Lemma 2.8 is taken without loss of generality as $e=0^{s}X0^{k}$, where $1\leq k\leq s$. ∎
Corollary 2.10.
There exists exactly one deleted path ${\cal P}(v)$ of length $s-k+1$ from the vertex $v=(0^{s}X0^{k-1})$, where $1\leq k\leq s$ and $X$ is a $s$-ones string.
Remark: Note that in this section we are considering trails and circuits in $G_{n-1}(s)$, but some of these trails, e.g., in the proof of Lemma 2.8 are paths and they are mentioned as such.
Lemma 2.11.
In $G_{n-1}(s)$, the first $s-i$ edges in a deleted path ${\cal P}(v)$ of $v=(0^{s}X_{1}0^{i})$ are disjoint from the first $s-j$ edges in a deleted path ${\cal P}(u)$ of $u=(0^{s}X_{2}0^{j})$, where $X_{1}$ and $X_{2}$ are two distinct $s$-ones strings and $i,j\geq 0$.
Proof.
The first $s-i$ edges in a deleted path that starts in the vertex $(0^{s}X_{1}0^{i})$ are exactly the edges either in the path
$0^{s}X_{1}0^{s}$
or in a path of the form
$0^{s}X_{1}0^{i+k}1Y,$
where $0\leq k<s-i$, and the length of $Y$ is $s-i-1-k$.
Similarly, the first $s-j$ edges in a deleted path that starts in the vertex $(0^{s}X_{2}0^{j})$ are exactly the edges either in the path
$0^{s}X_{2}0^{s}$
or in a path of the form
$0^{s}X_{2}0^{j+k}1Z,$
where $0\leq k<s-j$, and the length of $Z$ is $s-j-1-k$.
The binary representation of any vertex from the first $s$ vertices in these four paths starts in one of the zeros of the first run of $s$ zeros. Hence, a common vertex for these two paths implies that $X_{1}=X_{2}$, a contradiction. Thus, the claim of the lemma follows. ∎
Corollary 2.12.
The length of the circuit ${\cal C}$ (an $(n,s)$-sequence) in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces. If a circuit of such length exists it contains exactly all the words of the $(n,s)$-necklaces.
Proof.
Let $B$ be a necklace of order $n$ with a run of more than $s$ consecutive zeros. By Lemma 2.5 all the words in a necklace with a run of at least $12s+$ 1 zeros have at least one run with more than $s$ zeros and hence all the vertices of these necklaces are not contained in ${\cal C}$.
If the longest run of consecutive zeros in $B$ is $s+k$, $1\leq k\leq s$, then by Lemma 2.7 there are $n-1-s+k$ edges in the necklace with a run of more than $s$ consecutive zeros. Such a necklace has an edge $e$ in $G_{n-1}(s)$ of the form $e=0^{s}X0^{k}$, where $1\leq k\leq s$ and $X$ is an $s$-ones string, whose first vertex is $v=(0^{s}X0^{k-1})$ and by Lemma 2.8 it yields a deleted path ${\cal P}(v)$ whose length is at least $s-k+1$. Together, $B$ ($n-1-s+k$ forbidden edges) and the associated deleted path ${\cal P}(v)$ (at least $s-k+1$ forbidden edges) have at least $(n-1-s+k)+(s-k+1)=n$ edges which are not contained in ${\cal C}$. By Lemma 2.11 all the deleted paths of different vertices are disjoint and hence the number of edges of $G_{n-1}(s)$ which are not contained in ${\cal C}$ is at least the number of edges in the necklaces of order $n$ which contain edges with a run of more than $s$ consecutive zeros.
Thus, the length of the circuit ${\cal C}$ in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces in $G_{n-1}(s)$.
For the fact that such a circuit cannot contain any $(n,s)$-word not from an $(n,s)$-necklace we have to add the consequence of Corollary 2.10. ∎
What about the length of the maximum length trail ${\cal P}$ in $G_{n-1}(s)$? All the arguments we have used so far hold also for a trail with one exception. The trail can contain two vertices $v$ and $u$, where the in-degree of $v$ is one and its out-degree is two; the in-degree of $u$ is two and its out-degree is one. The first vertex in the trail is $v$ and the last vertex in the trail is $u$. We can apply the process in the proof of Lemma 2.8 on all the vertices in $G_{n-1}(s)$ except for $v$. After all the deleted paths are removed from $G_{n-1}(s)$ if we continue and apply the process in the proof of Lemma 2.8 on $v$ we will obtain a deleted path ${\cal P}(v)$ which starts in $v$ and ends at $u$. This deleted path can be added to the maximum length $(n,s)$-sequence to obtain a maximum length acyclic $(n,s)$-sequences. In other words, we have two vertices $v_{1}$ and $v_{2}$ for which we have the edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$ in $G_{n-1}(s)$ and the in-degree of $v$ is one. Such vertices have the form $v=(0^{s}X)$, $v_{1}=(0^{s-1}X0)$, and $v_{2}=(0^{s-1}X1)$. Note, that $v$ is a vertex contained in the $(n,s)$-necklace $[10^{s}X]$ and $v$ is also a vertex contained in the necklace $[0^{s+1}X]$ which is not an $(n,s)$-necklace, but $s$ words of the necklace are $(n,s)$-words. The maximum length circuit in $G_{n-1}$ cannot contain both edges $v\rightarrow v_{1}$ and $v\rightarrow v_{2}$ since the in-degree of $v$ is one. But, a trail can contain both edges if one of them will be the first edge in the trail. In other words, if both edges are on the trail ${\cal P}$, then since they have the same unique predecessor in $G_{n-1}(s)$, it follows that one of them must be the first edge in ${\cal P}$. It can be only at the beginning of a sub-path of ${\cal P}$ which was excluded by Lemma 2.8. The length of this sub-path in Lemma 2.8 is $s$ and its structure is as follows:
$(0^{s}X)\rightarrow(0^{s-1}X0)\rightarrow~{}\cdots~{}\rightarrow(0X0^{s-1})\rightarrow(X0^{s})~{},$
where all the $s$ edges are $(n,s)$-words that are contained in the necklace $[0^{s+1}X]$ which is not an $(n,s)$-necklace. Thus, we have the following conclusion.
Theorem 2.13.
The length of a trail ${\cal P}$ of maximum length in $G_{n-1}(s)$ is at most the number of edges in all the $(n,s)$-necklaces plus $s$ (unless the $(n,s)$-necklaces contain all the $(n,s)$-words in which case the bound on the length of a trail is the same as the one on the length of a circuit).
Example 3.
If $n=3$ and $s=2$, then an $(3,2)$-sequence of maximum length is
$0010111$
and an acyclic $(3,2)$-sequence of maximum length is
$001011100~{}.$
Both sequences have seven $(3,2)$-words and they contain all the $(3,2)$-necklaces.
If $n=5$ and $s=2$, then an $(5,2)$-sequence of maximum length has length 21 and one such sequence is.
001010011101011011111.001010011101011011111001010011101011011111~{}.001010011101011011111 .
An acyclic $(5,2)$-sequence has length 27 (4 additional bits for the acyclic representation and 2 additional bits from words which are not contained in $(5,2)$-necklaces) as follows
$0$
$\blacksquare$
A maximum length acyclic $(n,s)$-sequence (trail) in $G_{n-1}(s)$ can contain $s$ words which are not contained in a maximum length $(n,s)$-sequence. It was explained how to add them to a maximum length $(n,s)$-sequence. The maximum length trail can be constructed from an $(n,s)$-sequence in which one deleted path ${\cal P}(v)$ was taken from a necklace and it was added to obtain a maximum length $(n,s)$-sequence. It should be noted that the sequence starts and ends with the two possible shortest deleted paths. This is demonstrated in the following example.
Example 4.
The sequence
$00110101111101100101$
is a $(5,2)$-sequence of length 20 which contains two $(5,2)$-words $(00110)$ and $(01100)$ which are not contained in $(5,2)$-necklaces. This sequence falls short by one from the upper bound of a maximum length $(5,2)$-sequence. The deleted path which was removed from $G_{4}(2)$ is
$(0011)\rightarrow(0111)\rightarrow(1110)\rightarrow(1100)$
and its words of length 5 are contained in the $(5,2)$-necklace $[00111]$.
The sequence
$001101011111011001010011100$
of length 27 is an acyclic $(5,2)$-sequence which contains 23 $(5,2)$-words, two of which are not contained in an $(5,2)$-necklace. This sequence attains the upper bound of a maximum length acyclic $(5,2)$-sequence. In this sequence the deleted path was added at the end of the sequence and hence all the words of the $(5,2)$-necklaces are contained in this trail.
$\blacksquare$
3 Construction of Sequences of Maximum Length
In this section, we will concentrate on constructing $(n,s)$-sequences of maximum length. We will show that there exist $(n,s)$-sequences which attain the upper bound of Corollary 2.12 and hence these sequences are maximum length $(n,s)$-sequences. We will design a method and an efficient algorithm to generate a large class of maximum length $(n,s)$-sequences. The algorithms which will be designed are implemented for contructing the next bit given the last $n$ constructed bits of the sequence. We start by describing the general method, continue with an algorithm to construct a very large class of maximum length $(n,s)$-sequences, but with a large amount of storage and the complexity for computing the next bit of the sequence is not linear if $n$ is large. This algorithm is quite efficient when $n$ is not large. We finish with an algorithm that requires a more reasonable storage size, but with a slightly smaller number of generated sequences, but the best possible for the amount of storage that is used. The complexity for finding the next bit of the sequence is linear in $n$. The following method and its algorithm is a modification of the method and the algorithms used frequently to generate de Bruijn sequences, such as the one in [9]. While the upper bound on the length of a maximum length $(n,s)$-sequences was based on an analysis of trails in $G_{n-1}(s)$, the $(n,s)$-sequences which will be generated are cycles in $G_{n}(s)$. The basic principle in the method is taking all the $(n,s)$-necklaces and merging them into one cycle. For a vertex $X=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ in $G_{n}$, its companion $X^{\prime}$ is defined by $X^{\prime}=(x_{1},x_{2},\ldots,x_{n-1},\bar{x}_{n})$, where $\bar{x}$ is the binary complement of $x$. Two vertex-disjoint cycles in $G_{n}$, ${\cal C}_{1}$ and ${\cal C}_{2}$, are joined together into one cycle, if they contain a pair of companion states $X$ on ${\cal C}_{1}$ and $X^{\prime}$ on ${\cal C}_{2}$. If $Y\rightarrow X$ is an edge on ${\cal C}_{1}$ and $Z\rightarrow X^{\prime}$ is an edge on ${\cal C}_{2}$, then keeping all the edges of ${\cal C}_{1}$ and ${\cal C}_{2}$ except for these two edges and adding the edges $Y\rightarrow X^{\prime}$ and $Z\rightarrow X$, will merge the two cycles into one cycle. The merging of all the cycles is done as follows. The $(n,s)$-necklaces are ordered by their weight from the one with the largest weight to the one with the smallest weight. We start with the necklace which contains only ones and continue to the necklace with a unique zero and start with step 2. At each step, we have a main cycle composed of the $(n,s)$-necklaces that were merged so far and also the remaining $(n,s)$-necklaces. At step $i$, $i\geq 2$, we choose on each $(n,s)$-necklace ${\cal X}$ with weight $n-i$ (this is the number of ones in each word of the necklace) a vertex whose representation ends with a zero, e.g., $X0$. Its companion $X1$ on a necklace ${\cal X}^{\prime}$ has weight $n-i+1$ and it is readily verified that ${\cal X}^{\prime}$ does not have a run of more than $s$ zeros, i.e., ${\cal X}^{\prime}$ in also an $(n,s)$-necklace. Hence, $X1$ is on the main cycle, so we can merge the necklace ${\cal X}$ into the main cycles and continue. This procedure ends when all the $(n,s)$-necklaces are merged into the main cycle.
To apply this procedure, we have to choose on each $(n,s)$-necklace whose weight is at most $n-2$ a vertex whose binary representation ends with a zero. For example, this vertex can be the one which is the smallest one as a binary number among all the vertices whose binary representation ends with a zero. The companion of this vertex has a weight greater by one and hence its cycle was merged before and hence each necklace can be merged in its time to the main cycle. This procedure will result in a cycle of $G_{n}$ which contains all the vertices of all the $(n,s)$-necklaces. If we want to form more than one cycle we have to choose two vertices that end in a zero, e.g., one vertex ending in a zero with the smallest binary representation and a second one ending in a zero with the largest binary value. Some necklaces have only one word ending in a zero. On such a necklace ${\cal X}$ the unique word which ends with a zero has the form $(1^{d-1}0)^{n/d}$ which implies that the length $d$ of the necklace is a divisor of $n$. On these necklaces, we cannot have two choices. Fortunately, the number of such necklaces is less than the number of divisors of $n$ and we can ignore them. We should use a binary vector whose length is the number of $(n,s)$-necklaces. This vector indicates which one of the two vertices should be taken to perform the merging of the cycles. If $K$ is the number of $(n,s)$-necklaces of weight at most $n-2$ in which there are at least two words ending in a zero, then the algorithm can be used to generate $2^{K}$ distinct maximum length $(n,s)$-sequences. There is one disadvantage in this method – $K$ is exponential in $n$ and hence the algorithm will require an exponential area of storage. To overcome this disadvantage we can choose two such vertices only for some of the $(n,s)$-necklaces, for example, all the $(n,s)$-necklaces of some given weights.
Algorithm to merge $(n,s)$-necklaces:
Choose two integers $21\leq k_{1}\leq k_{2}\leq n-$ 2. For each $k$, $k_{1}\leq k\leq k_{2}$, use an order of the $(n,s)$-necklaces (with at least two words ending in a zero) whose words have weight $k$, using any given efficient algorithm. For each such $k$ choose a binary vector $V(k)$ whose length is the number of such $(n,s)$-necklaces with weight $k$. Using the order of these necklaces, the $i$th entry of $V(k)$ (associated with the $i$th necklace in this order) is set to 0 if we merge the $i$-necklace through the word of the smallest value and is set to 1 if we merge through the word with the largest value.
Initially, $\beta_{0}=(1,1,\ldots,1)=1^{n}$, $\beta_{1}=(1,1,\ldots,1,0)=1^{n-1}0$ and $w(\beta_{1})=n-1$. Given $\beta_{i}=(b_{i},b_{i+1},\ldots,b_{i+n-1})$ and its weight $w(\beta_{i})$, proceed to produce the next bit which implies the next $n$-tuple $\beta_{i+1}=(b_{i+1},\ldots,b_{i+n-1},b_{i+n})$ as follows.
  1. (A1)
    Let $\beta^{*}_{i}=(b_{i+1},\ldots,b_{i+n-1},0)$ and $w(\beta^{*}_{i})=w(\beta_{i})-b_{i}$ be the weight of $\beta^{*}_{i}$.
  2. (A2)
    If the necklace which contains $\beta^{*}_{i}$ contains a run of more than $s$ zeros, then go to (A5). If the necklace contains exactly one word ending in a zero, then go to (A6).
  3. (A3)
    If $k_{1}\leq w(\beta^{*}_{i})\leq k_{2}$, then let $d$ be the value of the entry in $V(w(\beta^{*}_{i}))$ associated with the necklace that contains $\beta^{*}_{i}$.
  4. (A4)
    Examine the cyclic shifts of $\beta^{*}_{i}$ for the existence of a shift $X$ which end with a zero whose value is the smallest one if $d=0$ (or if $w(\beta^{*}_{i})$ is not between $k_{1}$ and $k_{2}$) or whose value is the largest one if $d=1$. If $\beta^{*}_{i}=X$, then go to (A6).
  5. (A5)
    Set $b_{i+n}:=b_{i}$, $w(\beta_{i+1}):=w(\beta_{i})$, and stop.
  6. (A6)
    Set $b_{i+n}:=\bar{b}_{i}$, $w(\beta_{i+1}):=w(\beta_{i})-b_{i}+\bar{b}_{i}$.
The correctness of the algorithm and the proof of the next theorem is done similarly to the related proof in [9].
Theorem 3.1.
  1. (a)
    For every choice of $k_{1}$ and $k_{2}$, algorithm merge $(n,s)$-necklaces produces the next bit of a maximum length $(n,s)$-sequence.
  2. (b)
    For a given choice of $k_{1}$ and $k_{2}$, there are $2^{K}$ distinct choices of vertices for merging the $(n,s)$-necklaces, where $K$ is the number of $(n,s)$-necklaces with weight between $k_{1}$ and $k_{2}$ having at least two words which end in a zero. Thus, algorithm merge$(n,s)$-necklaces can be used to produce $2^{K}$ distinct maximum length $(n,s)$-sequences.
  3. (c)
    The working space that algorithm merge $(n,s)$-necklaces requires to produce the next bit of a maximum length $(n,s)$-sequence is $3n+K$ bits and the work required to produce the next bit is at most $n$ cyclic shifts and about the same number of $n$-bits comparisons.
Other algorithms for generating de Bruijn sequences can be also adapted to merge all the $(n,s)$-necklaces for constructing $(n,s)$-sequences. The order of the necklaces can be done by some known algorithms as the one by Kociumaka, Radoszewski, and Rytter [19]. But, if $n$ is large, then the known algorithms for ordering the necklaces, including the one in [19] are not linear and it will increase the time complexity of our overall algorithm. If $n$ is small, then we can use this method as the complexity will not be an issue.
To overcome the complexity problem, assuming that $n$ is large enough, we suggest choosing in most cases the word that ends in a zero with the maximal value for merging the necklace into the main cycle, but in some cases to use other words as follows.
Choose a large integer $k$ which depends on the number of maximum length $(n,s)$-sequences that we want to generate. On most $(n,s)$-necklaces the word the word ending in a zero with maximum value will be used to merge the necklace to the main cycle. On $k$ necklaces it will be another word. Each such a word is defined by $\theta(n,s,k)$ free parameters and hence $2^{k\cdot\theta(n,s,k)}$ maximum length $(n,s)$-sequences will be generated by the algorithm. The value of $\theta(n,s,k)$ will be increases as $n$ increases, or $s$ increases, or $k$ increases. Let $m=\lceil\log k\rceil$ and consider the ordered set $V=\{V(i)\}_{i=0}^{k-1}$ of $k$ distinct $(n,s)$-words in $k$ distinct $(n,s)$-necklaces, constructed as follows:
  1. (c1)
    The first $(s+1)\lceil\frac{m}{s}\rceil+1$ bits of $V(i)$ contain ones at positions $(s+1)j$, $0\leq j\leq\lceil\frac{m}{s}\rceil$ (this guarantee that in these position there will not be a run of more than $s$ consecutive zeros). In the other $s\lceil\frac{m}{s}\rceil$ positions (between the positions of the ones) $V(i)$ forms the binary representation of $i$. These $(s+1)\lceil\frac{m}{s}\rceil+1$ bits are followed by a single zero.
  2. (c2)
    The last $(s+1)\lceil\frac{m}{s}\rceil+4$ bits of $V(i)$ start and end with single zero and between these two zeros there are $(s+1)\lceil\frac{m}{s}\rceil+2$ ones. We want that this run of $(s+1)\lceil\frac{m}{s}\rceil+2$ ones will be the unique longest run of ones in the word.
  3. (c3)
    For each $j\geq 1$, in position $\left((s+1)\lceil\frac{m}{s}\rceil+2\right)j$, where $\left((s+1)\lceil\frac{m}{s}\rceil+2\right)j\leq n-(s+1)\lceil\frac{m}{s}\rceil-3$, there is a zero. This guarantee that the requirement for the longest run of ones in (c2) is satisfied.
  4. (c4)
    Between the first $(s+1)\lceil\frac{m}{s}\rceil+1$ bits and the last $(s+1)\lceil\frac{m}{s}\rceil+4$ bits each $s+1$ bits we have a one (at the end of these $s+1$ positions). If there is a collision between these ones and the zeros of (c3), then at the place of the collision we will have the one after $s$ bits and not after $s+1$ bits and after this one the zero (to avoid such a collision). This guarantees that there will not be a run of more than $s$ zeros in this section of the word.
  5. (c5)
    The rest of the positions which were not specified. These are free positions in which there are arbitrary assignments of zeros and ones.
We first note that each word of $V$ is an $(n,s)$-word from an $(n,s)$-necklace since the word starts with a one (by (c1)) and ends with an isolated zero (by (c2)) and in each $s+1$ consecutive positions there is at least one one by (c1), (c2), (c3), and (c4). Note further that each $V(i)$ has a unique run of exactly $(s+1)\left(\lceil\frac{m}{s}\rceil\right)+2$ ones which ends just one position before the last position . All the other runs of ones in each $V(i)$ are shorter. These properties are guaranteed by (c2) and (c3). This implies that two such words from $V$ cannot be a shift of the other and also all the words are of full-order. In the first $(s+1)\left(\lceil\frac{m}{s}\rceil\right)+1$ bits there is the binary representation of $i$ (by (c1)) with separations of ones to satisfy the constraint of no more than $s$ consecutive zeros (again by (c1)). This will enable to find if the word that we consider is from the set $V$. There are many free parameters whose number is increased as $s$ get larger and/or if we choose larger $k$. The algorithm will have similar steps to algorithm merge $(n,s)$-necklaces. The algorithm will first try to see if the necklace of $\beta_{i}^{*}$ is a word from $V$ in which case the necklace is joined to the main cycle only if $\beta_{i}^{*}$ is the associated word of $V$. If $\beta_{i}^{*}$ is not a word in the set $V$, then it examines whether $\beta_{i}^{*}$ is word ending in a zero with the maximum value in its necklace. Note, that since the largest run of ones is at the end of the word followed by a unique zero, it follows that no $V(i)$ can be the word of maximum value end in a zero in its necklace (this longest run of ones starts the word of the maximum value in the necklace. The procedure itself is very similar, but slightly more complicated, than the one described in [9].
The construction of the maximum length $(n,s)$-sequences is based on merging all the $(n,s)$-necklaces. There are $(n,s)$-sequences that contain $(n,s)$-words that are not contained in $(n,s)$-necklaces. But, these sequences cannot be of maximum length and this can be proved as a consequence from the proofs of Lemmas 2.8 and 2.12.
Finally, we consider a maximum length acyclic $(n,s)$-sequence. This is done by adding a short path with $(n,s)$-words, which are not contained in $(n,s)$-necklaces, to the maximum length $(n,s)$-sequence ${\cal C}$. Consider an $(n,s)$-word $(10^{s}X)$, for some $s$-ones string $X$, whose length is $n-1-s$, on an $(n,s)$-necklace with the edge $(10^{s}X)\rightarrow(0^{s}X1)$ of the associated necklace on the cycle ${\cal C}$. We start the maximum length path with the vertex $(10^{s}X)$ and continue with the edge $110^{s}X$ 1 to the vertex $(0^{s}X1)$ as on the cycle ${\cal C}$ until we reach the vertex $(10^{s}X)$ again. So far the path is the same as the cycle ${\cal C}$. Now, we continue with the $s$ edges,
$(10^{s}X)\rightarrow(0^{s}X0)\rightarrow~{}\cdots~{}\rightarrow(0X0^{s})~{},$
to obtain an acyclic $(n,s)$-sequence of maximum length, i.e., $s$ edges and $s$ vertices were added to the maximum length $(n,s)$-sequence ${\cal C}$ and the outcome is an acyclic $(n,s)$-sequence which attain the bound of Theorem 2.13. Note, that all the added vertices are on a necklace which is not an $(n,s)$-necklace, but the added vertices are represented by the $(n,s)$-words which are contained in the necklace. Note, further that the added sub-path is to the end of the cycle ${\cal C}$, while in the upper bound it was considered in its beginning, but this is equivalent.
4 Enumeration of the Number of Sequences
In this section, we will consider enumeration associated with the number of $(n,s)$-words, $(n,s)$-necklaces, and $(n,s)$-sequences related to the bounds and the construction that were presented in the previous sections. For this purpose, we define two important concepts, the rate and the redundancy of the set of sequences discussed in this paper. The rate of an $(n,s)$-sequence with maximum length $N$ will be defined as
${\cal R}(n,s)=\limsup_{n\rightarrow\infty}\frac{\log N}{n}~{}.$
The redundancy of an $(n,s)$-sequence ${\mathcal{S}}$ of length $N$ will be defined as
$\text{red}({\mathcal{S}})=n-\log N~{}.$
The algorithm which uses a de Bruijn sequence of length $2^{n}$ and each one is simulated by ’11’ and each zero by ’10’ considers a sequence of length $2^{n+1}$ with windows of length $2n$ for each required $n$-tuple. Hence, it has a high redundancy of $12n-\log 2^{n+1}=n-$ 1. Our $(n,s)$-sequences reduce this redundancy quite dramatically.
What is the exact length of a maximum length $(n,s)$-sequence? The computation of this size can be done similarly to other computations associated with constrained codes [15, 21], and asymptotic computations are done in the same way. Some of the computations can be done more accurately. For example, we will present the computations for $s=1$. Let $g_{n}$ be the number of $(n,1)$-words.
Lemma 4.1.
The number of $(n,1)$-words is
$g_{n}=\frac{\left(\frac{1+\sqrt{5}}{2}\right)^{n+2}-\left(\frac{1-\sqrt{5}}{2}\right)^{n+2}}{\sqrt{5}}~{}.$
Proof.
It is well-known and easily computed that $g_{n}=g_{n-1}+g_{n-2}$, where $g_{1}=2$ and $g_{2}=3$. The solution for this recurrence is $g_{n}=\frac{\varphi^{n+2}-\psi^{n+2}}{\sqrt{5}}$, where $\varphi=\frac{1+\sqrt{5}}{2}$ and $\psi=\frac{1-\sqrt{5}}{2}$. ∎
Lemma 4.2.
The number of $(n,1)$-words in the $(n,1)$-necklaces is $g_{n}-g_{n-4}$.
Proof.
A necklace that is not an $(n,1)$-necklace contains $(n,1)$-words if it starts with 01 and ends with 10. In between we have an $(n-4,1)$-word and hence by definition there are $g_{n-4}$ such words. Thus, the number of $(n,1)$-words in the $(n,1)$-necklaces is $g_{n}-g_{n-4}$. ∎
Now we can combine Corollary 2.12 and the constructions in Section 3 with Lemma 4.2 to obtain the following conclusion.
Corollary 4.3.
The length of a maximum length $(n,1)$-sequence is $g_{n}-g_{n-4}$.
Corollary 4.4.
The length of a maximum length acyclic $(n,1)$-sequence is $g_{n}-g_{n-4}+1$.
Corollary 4.5.
The rate of maximum length $(n,1)$ sequences is 0.6942 and their redundancy is $0.3058n$.
Lemma 4.1 can be simply generalized to $(n,s)$-words as follows.
Lemma 4.6.
The number $h_{n,s}$ of $(n,s)$-words satisfy the following recursive formula:
$h_{n,s}=\sum_{i=1}^{s+1}h_{n-i,s}~{},$
where $h_{i,s}=2^{i}$ for $1\leq i\leq s$ and $h_{s+1,s}=2^{s+1}-1$.
Proof.
Each $(n,s)$-word can start in $i$ zeros, $0\leq i\leq s$, followed by a one and after that there is an $(n-i-1,s)$-word. This implies the claim in the lemma. ∎
It was proved in [2, 20, 22] that $h_{n,s}$ can be expressed as
$h_{n,s}=\left\lfloor\frac{\lambda^{k+1}(\lambda-1)}{(s+2)\lambda-2(s+1)}+0.5\right\rfloor,$
where $\lambda$ is the unique positive real root of the equation
$x^{s+1}-\sum_{i=0}^{s}x^{i}=0~{}.$ (1)
Lemma 4.2 can be generalized as follows.
Lemma 4.7.
The number of $(n,s)$-words in the $(n,s)$-necklaces is
$h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i,s}~{}.$
Proof.
A necklace which is not an $(n,s)$-necklace contains $(n,s)$-words if such a word starts with $10^{i}$ 1 and ends with $10^{j}$, where $1\leq i,j\leq s$ and $s<i+j$. In between these two substrings we have an $(n-2-i-j,s)$-word and hence by definition, there are $h_{n-2-2s,s}$ such words if $i=j=s$. If $i=s-1$, $j=s$ or $i=s$, $j=s-1$, and $s>1$, then there are $2h_{n-1-2s,s}$ such words (for both together). If $i+j=2s-2$ and $s>2$, then $i=s-2$ or $i=s-1$ or $i=s$ and hence there are $3h_{n-2s,s}$ such words. This analysis is continued for each $s+1\leq i+j\leq 2s$, where $1\leq i,j\leq s$. Finally, if $i+j=s+1$, then $1\leq i\leq s$ and hence there are $s\cdot h_{n-s+3,s}$ such words. Thus, the number of $(n,s)$-words in the $(n,s)$-necklaces is $h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i,s}$. ∎
Corollary 4.8.
The length of a maximum length $(n,s)$-sequence is $h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i}$.
Corollary 4.9.
The length of a maximum length acyclic $(n,s)$-sequence is $h_{n,s}-\sum_{i=1}^{s}i\cdot h_{n-2s-3+i}+s$.
Corollary 4.10.
The rate of maximum length $(n,s)$ sequences is $\log\left|\lambda\right|$ and their redundancy is $n-\left|\lambda\right|n$, where $\lambda$ is the root of Eq. (1).
We are also interested in the number of generated cycles which is $2^{K}$, where $K$ is the number of $(n,s)$-necklaces whose weights are between $k_{1}$ and $k_{2}$. This is associated with the algorithm to merge $(n,s)$-necklaces. When $s=1$ this number is not difficult to compute.
Lemma 4.11.
The number of $(n,1)$-words with weight $k$ is $\binom{k+1}{n-k}$ if $k+1\geq n-k$.
Proof.
An $(n,1)$-word with weight $k$ has $k$ ones and the $n-k$ zeros are isolated between the ones (including at the beginning or the end) and hence there are $\binom{k+1}{n-k}$ such words. ∎
Lemma 4.12.
The number of words in the $(n,1)$-necklaces with words of weight $k$ is
$\binom{k}{n-k}+\binom{k-1}{n-k-1}$
if $k+1\geq n-k$.
Proof.
A word in an $(n,1)$-necklace is an $(n,1)$-word that does not start and end in a zero. By Lemma 4.11 the total number of $(n,1)$-words with weight $k$ is $\binom{k+1}{n-k}$. An $(n,1)$-word which starts with a zero and ends with a zero, starts with 01 and ends with 10. In between we have an $(n-4,1)$-word with weight $k-2$. By Lemma 4.11, there are $\binom{k-1}{n-k-2}$ such words and hence the number of words in the $(n,1)$-necklaces with words of weight $k$ is
$\binom{k+1}{n-k}-\binom{k-1}{n-k-2}=\binom{k}{n-k}+\binom{k-1}{n-k-1}~{}.$
It should be noted that on one hand the set $V$ defined in Section 3 is less effective when $s=1$, but based on Lemmas 4.11 and 4.12 an efficient algorithm to construct a large set of maximum length $(n,1)$-sequences can be designed. This algorithm will be based on algorithm merge $(n,1)$-necklaces and an efficient algorithm to enumerate the associated binomial coefficient, e.g., the enumerative encoding of Cover [5].
5 Sequences with Efficient Positioning Decoding
[…]
For ordering the necklaces, it is worthwhile to know their number. The total number of necklaces of order $n$, $Z(n)$, is a well-known result [8, 13]. It involves the well-known arithmetic function, the Möbius function [17].
Theorem 5.1.
The number of necklaces of order $n$ is
$Z(n)=\frac{1}{n}\sum_{d|n}\mu(d)2^{n/d},$
where $\mu(\cdot)$ is the Möbius function.
The Lyndon words have some simple properties.
Lemma 5.2.
The Lyndon word of a necklace starts with the longest run of zeros and ends with a one.
Lemma 5.3.
If the Lyndon word $X_{1}$ of a necklace has a larger run of consecutive zeros than in the Lyndon word $X_{2}$ of another necklace, than $x_{1}$ has a smaller value than $X_{2}$.
Lemma 5.2 is not a precise characterization of a Lyndon word since for example there might be a few runs with the longest run of zeros.
The least lexicographic de Bruijn sequence is generated based on the following celebrated lemma [12].
Lemma 5.4.
Let $X_{1},X_{2},X_{3},~{}\cdots,~{}X_{Z(n)}$ be the ordering of the Lyndon words of order $n$ from the smallest one lexicographically to the largest one. The concatenation of these words in this order is a de Bruijn sequence of order $n$, i.e., an Eulerian circuit in $G_{n-1}$ which is also a Hamiltonian cycle in $G_{n}$.
The next step is to show that lemma 5.4 is also true if we restrict ourselves to the Lyndon words of the $(n,s)$-necklaces.
Lemma 5.5.
Let $X_{1},X_{2},X_{3},~{}\cdots,~{}X_{Z(n)}$ be the ordering of the Lyndon words of order $n$ from the smallest one lexicographically to the largest one. There exists an index $k$ such that $X_{1},X_{2},~{}\cdots,~{}X_{k}$ are Lyndon words which are not contained in $(n,s)$-necklaces, while $X_{k+1},X_{k+2},~{}\cdots,~{}X_{Z(n)}$ are Lyndon words which are contained in $(n,s)$-necklaces.
Proof.
Since the Lyndon words are ordered by their lexicographic value from the smallest to the largest, it follows that they are first ordered by the number of leading zeros (from the largest number to the smallest number of leading zeros) in each word. By Lemma 5.2, this number of leading zeros is also the longest run of zeros in the necklace. By Lemma 5.3, all Lyndon words that start with more than $s$ zeros are contained in necklaces that are not $(n,s)$-necklaces, while all the Lyndon words that start with at most $s$ zeros are contained in $(n,s)$-necklaces which imply the claim of the lemma. ∎
Corollary 5.6.
The last Lyndon word, which is not an $(n,s)$-word, in the order from the smallest one to the largest one, is $0^{s+1}1^{n-s-1}$.
Proof.
By Lemma 5.3, the Lyndon words with a smaller number of leading zeros have larger values than the ones with a larger number of leading zeros. The word $0^{s+1}1^{n-s-1}$ is a Lyndon word with the largest value among the words starting with $s+1$ zeros. The next Lyndon word in the order has $s$ leading zeros and the claim of the lemma follows. ∎
Corollary 5.7.
In the lexicographic ordering of the Lyndon words of order $n$, all Lyndon words after the Lyndon word $0^{s+1}1^{n-s-1}$ are contained in $(n,s)$-necklaces.
Proof.
All Lyndon words after $0^{s+1}1^{n-s-1}$ start with at most $s$ zeros and end with a one and hence their necklaces cannot contain any run of more than $s$ zeros. This implies that all these words are contained in $(n,s)$-necklaces. ∎
The algorithm which generates all the Lyndon words that follow the last Lyndon word $0^{s+1}1^{n-s-1}$ which is not contained in an $(n,s)$-necklace is given as follows.
Algorithm for the lexicographic generation of Lyndon words in $(n,s)$-necklaces:
The first word is defined as $X_{0}=0^{s+1}1^{n-s-1}$ (this word is not considered for the $(n,s)$-necklaces); set $Y:=X_{0}$ and $i:=0$.
  1. (E1)
    If $Y=1^{n}$, goto (E5); otherwise set $i:=i+1$.
  2. (E2)
    Find the largest index $j$ in $Y=y_{1}y_{2}~{}\cdots~{}y_{n}$ such that $y_{j}=0$.
  3. (E3)
    Let $Z=y_{1}y_{2}~{}\cdots~{}y_{j-1}1$ and set $Y$ to be the first $n$ bits of the infinite sequence $ZZZ~{}\cdots$.
  4. (E4)
    If $j$ divides $n$, then set $X_{i}:=Z$ and goto (E1); otherwise goto (E2).
  5. (E5)
    $X_{1},X_{2},\ldots,X_{i}$ is the sequence of Lyndon words of the $(n,s)$-necklaces from the smallest to the largest.
The concatenation of $X_{1}X_{2}~{}\cdots~{}X_{i}$ is the required maximum length $(n,s)$-sequence that contains exactly all the words of the $(n,s)$-necklaces.
The correctness of the algorithm is shown in the same way as it is proved in the FKM algorithm. To find the position of a given word $v$ of length $n$ in the concatenation of the Lyndon words we have to apply the algorithm proposed by Kociumaka, Radoszewski, and Rytter [19] which finds the position of $v$ in the concatenation of all the Lyndon word of order $n$. After the position of $v$ was found we have to subtract the position of the last entry in the word $0^{s+1}1^{n-s-1}$ since our sequence starts after this last entry. It is worthwhile and more efficient in the long run to compute this position in advance and save it as it will be used in every application of the algorithm. The complexity of the decoding algorithm is the same as in [19] since the ordering of the Lyndon words is the same with the exception that we do not start from the first one but with the one after $0^{s+1}1^{n-s-1}$.
6 Conclusion and Discussion
Motivated by an application for space-free quantum key distribution a system based on a simple run-length limited sequences in the de Bruijn graphs is proposed. The maximum length of such sequences is shown to be associated with the number of constrained necklaces. An efficient algorithm to generate a large set of such sequences is proposed and some enumerations related to the length of a maximum length sequences are discussed. Another algorithm that can find efficiently the position of each word of length $n$ in the sequence is also presented.
The exposition of the previous sections can be generalized in a few directions. First, we can consider a non-binary alphabet with the same constraint on the number of zeros, i.e., the longest run of zeros has length at most $s$. The maximum length sequence will be of length which is the total number of words in the necklaces that do not contain a cyclic run of zeros whose length is larger than $s$. A general construction for a large set of maximum length $(n,s)$-sequences can be done similarly as was done in Section 3 using a generalization as done in [7]. Also a maximum length $(n,s)$-sequence in which the position of each word of length $n$ can be efficiently computed is done similarly to the binary case [19].
A different direction is to add more constraints on the sequences. For example, in addition to the requirement that each window of length $n$ has a run of at most $s$ consecutive zeros, we can require that the each window of length $n$ will contain at least $w$ ones.
An $(n,w,s)$-word is an $(n,s)$-word whose weight at least $w$. An $(n,w,s)$-necklace is a necklace for which all its vertices are $(n,w,s)$-words. Let $G_{n}(w,s)$ be the subgraph of $G_{n}$ induced by all the vertices of $G_{n}$ whose representations are $(n,w,s)$-words. The following theorem is proved with the same steps used in Corollary 2.12.
Theorem 6.1.
The length of the circuit ${\cal C}$ in $G_{n-1}(w,s)$ is at most the number of edges in all the $(n,w,s)$-necklaces.
We can merge all the $(n,w,s)$-necklaces with an algorithm which is similar to the algorithm for merging the $(n,s)$-necklaces with adaptation for the threshold on the weight of the windows of length $n$. Theorem 3.1 holds also for this algorithm.
We can also design a more efficient algorithm using a set of words on which the merging of the cycles will be done similarly as was done in the algorithm to merge all the $(n,s)$-necklaces. The ideas will be similar to the ones in [7] with the modification and adaption as was done in Section 3.
The constructed sequences have the potential to be used in other modern applications. A large set of sequences where each sequence can be encoded and decoded efficiently is a major task for the future. Moreover, these algorithms should be also used where the set of constrained words of length $n$ are mapped to a set of information words with a smaller length $k$.